# ExPairT-LLM: Exact Learning for LLM Code Selection by Pairwise Queries

### v1

@@ Abstract
Despite recent advances in LLMs, the task of code generation is still challenging. To cope, code selection algorithms select the best program from multiple programs generated by an LLM. However, existing algorithms can fail to identify the correct program, either because they can misidentify nonequivalent programs or because they rely on an LLM and assume it always correctly determines the output for every input. We present ExPairT-LLM, an exact learning algorithm for code selection that selects a program by posing to an LLM oracle two new types of queries: pairwise membership and pairwise equivalence. These queries are simpler for LLMs and enable ExPairT-LLM to identify the correct program through a tournament, which is robust to some LLM mistakes. We evaluate ExPairT-LLM on four popular code datasets. Its pass@1 (success rate) outperforms the state-of-the-art code selection algorithm on average by $13.0\%$ and up to $27.1\%$. It also improves the pass@1 of LLMs performing complex reasoning by $24.0\%$.
@@
Extended versionhttps://arxiv.org/abs/2511.xxxxx
@@ Introduction
Large language models (LLMs) show remarkable performance in code generation (Code_Llama; HumanEval; NijkampPHTWZSX23; LiAZMKMMALCLZZW23; DeepSeekCoder; LuoX0SGHT0LJ24; FengGTDFGS0LJZ20; Qwen2Coder). However, they may return code solutions that do not meet the task description, do not align with the input-output examples, or do not even compile (LiuXW023; Shihan24). Code selection algorithms let the LLM generate multiple programs and automatically identify the correct program among them. Some of these approaches rely on input-output examples generated by the LLM (ChenZNZLLC23; Mouxiang24; ToNB24); however, the outputs may be incorrect. Other works generate inputs, group programs based on their outputs, and select the cluster containing the most programs (alpahCode; ShiFGZW22). Another approach is to use neural networks to estimate the correctness of a program without relying on test inputs (InalaWYCELMG22; ZhangYHLYF023). However, these approaches can fail to identify the correct program if the input-output examples are incorrect or if the examples do not differentiate between nonequivalent programs, which can lead to returning an incorrect program. An exception is a work that looks for differentiating inputs through fuzzing and symbolic execution (FanRMR24). However, it is not robust to LLM errors, and fuzzing and symbolic execution tend to fail on tasks with complex input constraints (FanRMR24).
@@
Several code selection algorithms rely on LLMs to generate input-output examples, and then cluster programs by their consistency with the examples (ChenZNZLLC23; Mouxiang24; ToNB24). Others generate inputs, group programs by their outputs, and select the cluster with the most programs (alpahCode; ShiFGZW22). However, they can fail for various reasons: if the input-output examples are incorrect, if the selected cluster is incorrect, or if the selected cluster contains nonequivalent programs. LLMCodeChoice leverages fuzzing and symbolic execution to find inputs that differentiate the candidate programs (FanRMR24); however, it is unsuitable for tasks with complex input constraints and assumes the LLM is always correct. Others use neural networks to estimate the correctness of a program for a task without executing the program on test inputs (InalaWYCELMG22; ZhangYHLYF023).
@@
Relying on LLMs as oracles (judges) for selecting between two candidates has been proposed for text-related tasks, including summarization (LiusieMG24; LiusieRFG24; Yinhong24), chat assistants (LiusieMG24; LiusieRFG24; ZhengC00WZL0LXZ23), content generation (Yinhong24), and information retrieval (QinJHZWYSLLMWB24). CodeRL (Le0GSH22) employs reinforcement learning to enhance code-generating LLMs during training and inference by utilizing a separate LLM trained to assess the functional correctness of generated code samples. Other works utilize LLMs for self-repair in code generation models, providing feedback that identifies errors from execution results and generates explanations to fix the code (ChenLSZ24; ZhangLLLJ23; OlaussonIW0S24).
@@
A coding task consists of a natural language description (e.g., a docstring describing the function) and possibly input-output examples. Figure 2 shows task examples. If a program $p$ satisfies a task $t$, we say $p$ is correct and write $p\models t$. Given a task, we are given a finite set of candidate programs $\mathcal{P}$. We assume that there exists $p\in\mathcal{P}$ that is correct and that the programs are deterministic and terminate for every input (stochastic or nonterminating programs can be removed in a preprocessing step by running them on several inputs with a reasonable timeout). Obtaining the set of candidates is orthogonal to our problem.
@@
Given a task $t$ and a set of programs $\mathcal{P}$, containing a correct program, we design an algorithm $\mathcal{A}$ returning $p\in\mathcal{P}$ satisfying $t$. We evaluate a code selection algorithm $\mathcal{A}$ by its pass@1 (HumanEval) (success rate). Given a set of task-programs pairs $\mathcal{T}$, the pass@1 of $\mathcal{A}$ is the percentage of tasks for which it returns a correct program:
@@
This problem is challenging. First, tasks are described by natural language and examples, which may be ambiguous, under-specified, or may require domain knowledge, e.g., mathematics (Figure 2(a)). Second, the candidate programs are often too complex (Shihan24) to be analyzed by existing program analyzers, especially if the tasks pose complex input constraints (e.g., Figure 2(b)), which are not easily expressed with existing tools. Third, the algorithm cannot interact with a user to obtain more information. Our problem is not a classical program synthesis problem (GulwaniPS17; AlurBJMRSSSTU13). First, the program space can be complex, consisting of hundreds of lines of code, spanning a wide range of operations, data structures, and sometimes uncommon libraries. Second, the task may include a lengthy natural language description that is difficult to translate into a formal specification. Third, the tasks can vary significantly, e.g., mathematics (Figure 2(a)) and algorithms and data structures (Figure 2(b)), whereas program synthesizers often focus on related tasks.
@@
LLMs have been shown to be effective as judges for selecting between two candidates in text-related tasks (LiusieMG24; LiusieRFG24; Yinhong24; QinJHZWYSLLMWB24; ZhengC00WZL0LXZ23). Additionally, several works rely on pairwise comparisons to identify the best candidate from a set of candidates (WauthierJJ13; JamiesonN11; ShahW17; Cattelan12; Newman22). One robust approach is Copeland’s method (Cop51), a voting system that determines the winner by head-to-head comparisons between each pair of candidates. In every comparison, one candidate gains a point. The winner is the candidate with the most points. Theoretically, we could rely on Copeland’s method to identify a correct program by posing to the LLM oracle a series of queries, each presents two programs and asks the oracle to select the better program for the given task. However, such pairwise queries are still challenging for LLMs.
@@
Algorithm 1 shows its pseudocode. The inputs are a coding task in natural language $t$, a list of inputs $\mathcal{I}$ (provided by the user or LLM-generated), a list of candidate programs $\mathcal{P}$, and an LLM $\mathcal{L}$. It returns a program $p\in\mathcal{P}$, intended to satisfy $t$. It first initializes the selected cluster ${C}^{*}$ to $\mathcal{P}$ (Line 3). Next, it iteratively refines ${C}^{*}$ until ${C}^{*}$ contains only equivalent programs (Line 4). Each iteration begins by invoking the cluster function (Line 5). This function partitions ${C}^{*}$ into clusters $\mathcal{C}$ and their respective outputs $\mathcal{O}$, based on the outputs of programs in ${C}^{*}$ for the inputs $\mathcal{I}$ (Line 22–Line 31). Then, ExPairT-LLM initializes the scores array, which records the clusters’ scores (Line 6). Then, Copeland’s method is executed by posing a pairwise membership query for each pair of clusters (Line 7–Line 11). The cluster with the maximal score is set to $C^{*}$ (12). Next, ExPairT-LLM checks if the programs in $C^{*}$ are equivalent. It initializes the eq flag to True (Line 13). It then poses pairwise equivalence queries to the LLM, between the first program in $C^{*}$ and every other program in ${C}^{*}$ (Line 14–Line 15). If the LLM finds a differentiating input x for two programs, ExPairT-LLM validates it by executing both programs on x (Line 16). If the validation succeeds, ExPairT-LLM sets the inputs $\mathcal{I}$ to x, negates the flag eq, and begins another iteration to refine $C^{*}$ (Line 17). Otherwise, if eq remains True, ExPairT-LLM returns the first program in $C^{*}$ (Line 18). The query prompts are in the extended version.
@@
We next present theoretical guarantees and proof sketches. Full proofs are in the extended version. We recall that the program space $\mathcal{P}$ is finite, contains a correct program, and consists of terminating and deterministic programs. We first show that ExPairT-LLM terminates. Second, if the oracle $\mathcal{L}$ is accurate, ExPairT-LLM returns the correct program. Otherwise, we show a lower bound on the probability of it choosing the correct cluster and show the probability of identifying that a cluster contains nonequivalent programs. Lastly, we analyze the number of queries.
@@
We implemented ExPairT-LLM in Python. Experiments ran on an Ubuntu 20.04.6 OS on a dual AMD EPYC 7742 server with 1TB RAM. Unless otherwise specified, each task has $|\mathcal{P}|=25$ candidate programs, an initial input list $\mathcal{I}$ with five inputs generated by GPT-4o (GPT4), and OpenAI o1-mini (openaio1mini) as the LLM oracle. We evaluate on three datasets: HumanEval (HumanEval), which has 164 coding tasks; MBPP-sanitized (Austin21) (denoted as MBPP), which has 427 coding tasks; and APPS (hendrycks2021measuring), whose test set has 5,000 coding tasks. To fairly compare with prior works (Mouxiang24; ChenZNZLLC23), for HumanEval, we exclude input-output examples provided within the tasks. These datasets are the most popular benchmarks for evaluating code generation models (Code_Llama; DeepSeekCoder; Qwen2Coder). Each of these datasets provides test input-output examples for each task (hidden during the selection), which we use to evaluate the correctness of the selected programs. For each dataset, we consider several models for generating the program space $\mathcal{P}$: Codex (HumanEval) (code-davinci-002 version), CodeLlama (Code_Llama) (7B-Python version), StarCoder (LiAZMKMMALCLZZW23), DeepSeek-Coder (DeepSeekCoder) (6.7B Instruct version), and CodeGen (NijkampPHTWZSX23) (MONO-16.1B version). Due to the substantially larger number of coding tasks in APPS, we follow previous works (Mouxiang24; ChenZNZLLC23) and consider for its evaluation only the Codex model. The program space generated by each model is obtained from the repository provided by Mouxiang24. We evaluate a code selection algorithm by the popular pass@1 metric (Equation 1). Like Mouxiang24; ChenZNZLLC23, we exclude from the pass@1 pairs $(t,\mathcal{P})$ where all programs in $\mathcal{P}$ satisfy $t$, or where no program in $\mathcal{P}$ satisfies $t$.
@@
We compare ExPairT-LLM with two code selection algorithms. First, $\mathcal{B}^{4}$ (Mouxiang24), the state-of-the-art, which defines the optimal selection strategy using a Bayesian framework, where the posterior probability of passing states between solutions and tests guides the selection. It approximates this posterior using Bayesian statistics and reformulates the problem as integer programming. $\mathcal{B}^{4}$ has three variants. In our comparison, for each dataset and model, we define its pass@1 by the best variant. Second, CodeT (ChenZNZLLC23), which generates test cases for each task, executes the candidate programs on these cases, and selects a program based on dual execution agreement. This agreement considers both the consistency of the outputs with the generated test cases and their agreement with the outputs of other candidate programs. For both $\mathcal{B}^{4}$ and CodeT, we use the authors’ code, with the generated input-output examples provided in the repository of Mouxiang24. We also compare with the Original baseline, i.e., the LLM which returns a program given a task. Table 1 shows the results. ExPairT-LLM outperforms all baselines across all datasets and models. On average, its pass@1 exceeds $\mathcal{B}^{4}$ by $13.0\%$, CodeT by $16.6\%$, and Original by $40.7\%$. We assess the statistical significance of these improvements with McNemar’s test (McNemar47), suitable for comparing two algorithms on the same set of tasks with a binary success metric (e.g., pass@1). It confirms that ExPairT-LLM significantly outperforms each baseline on every dataset ($p$-value <0.001<0.001).

### v2

@@ Abstract
Despite recent advances in LLMs, the task of code generation is still challenging. To cope, code selection algorithms select the best program from multiple programs generated by an LLM. However, existing algorithms can fail to identify the correct program, either because they fail to distinguish nonequivalent programs or because they rely on an LLM and assume it always correctly determines the output for every input. We present ExPairT-LLM, an exact learning algorithm for code selection that selects a program by posing two new types of queries to an LLM oracle: pairwise membership and pairwise equivalence. These queries are simpler for LLMs and enable ExPairT-LLM to identify the correct program through a tournament, which is robust to some LLM mistakes. We evaluate ExPairT-LLM on four popular code datasets. Its pass@1 (success rate) outperforms the state-of-the-art code selection algorithm on average by $13.0\%$ and up to $27.1\%$. It also improves the pass@1 of LLMs performing complex reasoning by $24.0\%$.
@@
Extended versionhttps://arxiv.org/abs/2511.10855
@@ Introduction
Large language models (LLMs) show remarkable performance in code generation (Code_Llama; HumanEval; NijkampPHTWZSX23; LiAZMKMMALCLZZW23; DeepSeekCoder; LuoX0SGHT0LJ24; FengGTDFGS0LJZ20; Qwen2Coder). However, they may return code solutions that do not meet the task description, do not align with the input-output examples, or do not even compile (LiuXW023; Shihan24). Code selection algorithms let the LLM generate multiple programs and automatically identify the correct program among them. Some of these approaches rely on input-output examples generated by the LLM (ChenZNZLLC23; Mouxiang24; ToNB24); however, the outputs may be incorrect. Other works generate inputs, group programs based on their outputs, and select the cluster containing the most programs (alpahCode; ShiFGZW22). Another approach is to use neural networks to estimate the correctness of a program without relying on test inputs (InalaWYCELMG22; ZhangYHLYF023). However, these approaches can fail to identify the correct program if the input-output examples are incorrect or if the examples do not differentiate between nonequivalent programs, which can lead to returning an incorrect program. An exception is a method that looks for differentiating inputs through fuzzing and symbolic execution (FanRMR24). However, it is not robust to LLM errors, and fuzzing and symbolic execution tend to fail on tasks with complex input constraints (FanRMR24).
@@
Several code selection algorithms rely on LLMs to generate input-output examples and cluster programs by their consistency with the examples (ChenZNZLLC23; Mouxiang24; ToNB24). Other approaches generate inputs, group programs by their outputs, and select the cluster with the most programs (alpahCode; ShiFGZW22). However, they can fail for various reasons: if the input-output examples are incorrect, if the selected cluster is incorrect, or if the selected cluster contains nonequivalent programs. LLMCodeChoice uses fuzzing and symbolic execution to find inputs that differentiate candidate programs (FanRMR24); however, it is unsuitable for tasks with complex input constraints and assumes the LLM is always correct. Others use neural networks to estimate the correctness of a program for a task without executing the program on test inputs (InalaWYCELMG22; ZhangYHLYF023).
@@
Relying on LLMs as oracles (judges) for selection between two candidates has been proposed for text-related tasks, including summarization (LiusieMG24; LiusieRFG24; Yinhong24), chat assistants (LiusieMG24; LiusieRFG24; ZhengC00WZL0LXZ23), content generation (Yinhong24), and information retrieval (QinJHZWYSLLMWB24). CodeRL (Le0GSH22) employs reinforcement learning to enhance code-generating LLMs during training and inference by utilizing a separate LLM trained to assess the functional correctness of generated code samples. Other works utilize LLMs for self-repair in code generation models, providing feedback that identifies errors from execution results and generates explanations to fix the code (ChenLSZ24; ZhangLLLJ23; OlaussonIW0S24).
@@
A coding task consists of a natural language description (e.g., a docstring describing the function) and possibly input-output examples. Figure 2 shows task examples. If a program $p$ satisfies a task $t$, we say $p$ is correct and write $p\models t$. For each task, we are given a finite set of candidate programs $\mathcal{P}$. We assume that there exists $p\in\mathcal{P}$ that is correct and that the programs are deterministic and terminate for every input (stochastic or nonterminating programs can be removed in a preprocessing step by running them on several inputs with a reasonable timeout). Obtaining the set of candidates is orthogonal to our problem.
@@
Given a task $t$ and a set of programs $\mathcal{P}$ that contains a correct program, a code selection algorithm $\mathcal{A}$ aims to return $p\in\mathcal{P}$ satisfying $t$. We evaluate algorithm $\mathcal{A}$ by its pass@1 (HumanEval) (success rate). Given a set of task-program-set pairs $\mathcal{T}$, the pass@1 of $\mathcal{A}$ is the percentage of tasks for which it returns a correct program:
@@
This problem is challenging. First, tasks are described by natural language and examples, which may be ambiguous, under-specified, or require domain knowledge, e.g., mathematics (Figure 2(a)). Second, the candidate programs are often too complex (Shihan24) to be analyzed by existing program analyzers, especially if the tasks pose complex input constraints (e.g., Figure 2(b)), which are not easily expressed with existing tools. Third, the algorithm cannot interact with a user to obtain more information. Our problem is not a classical program synthesis problem (GulwaniPS17; AlurBJMRSSSTU13). First, the program space can be complex, consisting of hundreds of lines of code, spanning a wide range of operations, data structures, and sometimes uncommon libraries. Second, the task may include a lengthy natural language description that is difficult to translate into a formal specification. Third, the tasks can vary significantly, e.g., mathematics (Figure 2(a)) and algorithms and data structures (Figure 2(b)), whereas program synthesizers often focus on related tasks.
@@
LLMs have been shown to be effective as judges for selecting between two candidates in text-related tasks (LiusieMG24; LiusieRFG24; Yinhong24; QinJHZWYSLLMWB24; ZhengC00WZL0LXZ23). Additionally, several works rely on pairwise comparisons to identify the best candidate from a set of candidates (WauthierJJ13; JamiesonN11; ShahW17; Cattelan12; Newman22). One robust approach is Copeland’s method (Cop51), a voting system that determines the winner by head-to-head comparisons between each pair of candidates. In every comparison, one candidate gains a point. The winner is the candidate with the most points. Theoretically, we could rely on Copeland’s method to identify a correct program by posing to the LLM oracle a series of queries, each of which presents two programs and asks the oracle to select the better program for the given task. However, such pairwise queries are still challenging for LLMs.
@@
Algorithm 1 shows its pseudocode. The inputs are a coding task in natural language $t$, a list of inputs $\mathcal{I}$ (provided by the user or LLM-generated), a list of candidate programs $\mathcal{P}$, and an LLM $\mathcal{L}$. It returns a program $p\in\mathcal{P}$, intended to satisfy $t$. It first initializes the selected cluster ${C}^{*}$ to $\mathcal{P}$ (Line 3). Next, it iteratively refines ${C}^{*}$ until ${C}^{*}$ contains only equivalent programs (Line 4). Each iteration begins by invoking the cluster function (Line 5). This function partitions ${C}^{*}$ into clusters $\mathcal{C}$ and their respective outputs $\mathcal{O}$, based on the outputs of programs in ${C}^{*}$ for the inputs $\mathcal{I}$ (Line 22–Line 31). Then, ExPairT-LLM initializes the scores array, which records the clusters’ scores (Line 6). Then, Copeland’s method is executed by posing a pairwise membership query for each pair of clusters (Line 7–Line 11). The cluster with the maximal score is set to $C^{*}$ (Line 12). Next, ExPairT-LLM checks if the programs in $C^{*}$ are equivalent. It initializes the eq flag to True (Line 13). It then poses pairwise equivalence queries to the LLM between the first program in $C^{*}$ and every other program in ${C}^{*}$ (Line 14–Line 15). If the LLM finds a differentiating input x for two programs, ExPairT-LLM validates it by executing both programs on x (Line 16). If the validation succeeds, ExPairT-LLM sets the inputs $\mathcal{I}$ to [x], negates the flag eq, and begins another iteration to refine $C^{*}$ (Line 17). Otherwise, if eq remains True, ExPairT-LLM returns the first program in $C^{*}$ (Line 18). The appendix show the query prompts.
@@
We next present theoretical guarantees and proof sketches. Full proofs are in the appendix. Recall that the program space $\mathcal{P}$ is finite, contains a correct program, and consists of terminating and deterministic programs. We first show that ExPairT-LLM terminates. Second, if the oracle $\mathcal{L}$ is accurate, ExPairT-LLM returns the correct program. Otherwise, we show a lower bound on the probability of it choosing the correct cluster and show the probability of identifying that a cluster contains nonequivalent programs. Lastly, we analyze the number of queries.
@@
We implemented ExPairT-LLM in Python. The experiments were run on an Ubuntu 20.04.6 OS on a dual AMD EPYC 7742 server with 1TB RAM. Unless otherwise specified, each task has $|\mathcal{P}|=25$ candidate programs, an initial input list $\mathcal{I}$ with five inputs generated by GPT-4o (GPT4), and OpenAI o1-mini (openaio1mini) as the LLM oracle. We evaluate on three datasets: HumanEval (HumanEval), which has 164 coding tasks; MBPP-sanitized (Austin21) (denoted as MBPP), which has 427 coding tasks; and APPS (hendrycks2021measuring), whose test set has 5,000 coding tasks. To fairly compare with prior works (Mouxiang24; ChenZNZLLC23), for HumanEval, we exclude input-output examples provided within the tasks. These datasets are the most popular benchmarks for evaluating code generation models (Code_Llama; DeepSeekCoder; Qwen2Coder). Each of these datasets provides test input-output examples for each task (hidden during the selection), which we use to evaluate the correctness of the selected programs. For each dataset, we consider several models for generating the program space $\mathcal{P}$: Codex (HumanEval) (code-davinci-002 version), CodeLlama (Code_Llama) (7B-Python version), StarCoder (LiAZMKMMALCLZZW23), DeepSeek-Coder (DeepSeekCoder) (6.7B Instruct version), and CodeGen (NijkampPHTWZSX23) (MONO-16.1B version). Due to the substantially larger number of coding tasks in APPS, we follow previous works (Mouxiang24; ChenZNZLLC23) and consider for its evaluation only the Codex model. The program space generated by each model is obtained from the repository provided by Mouxiang24. We evaluate a code selection algorithm by the popular pass@1 metric (Equation 1). Like Mouxiang24; ChenZNZLLC23, we exclude from the pass@1 pairs $(t,\mathcal{P})$ where all programs in $\mathcal{P}$ satisfy $t$, or where no program in $\mathcal{P}$ satisfies $t$.
@@
We compare ExPairT-LLM with two code selection algorithms. First, $\mathcal{B}^{4}$ (Mouxiang24), the state-of-the-art, defines the optimal selection strategy using a Bayesian framework, where the posterior probability of passing states between solutions and tests guides the selection. It approximates this posterior using Bayesian statistics and reformulates the problem as integer programming. $\mathcal{B}^{4}$ has three variants. In our comparison, for each dataset and model, we define its pass@1 by the best variant. Second, CodeT (ChenZNZLLC23), which generates test cases for each task, executes the candidate programs on these cases, and selects a program based on dual execution agreement. This agreement considers both the consistency of the outputs with the generated test cases and their agreement with the outputs of other candidate programs. For both $\mathcal{B}^{4}$ and CodeT, we use the authors’ code, with the generated input-output examples provided in the repository of Mouxiang24. We also compare with the Original baseline, i.e., the LLM which returns a program given a task. Table 1 shows the results. ExPairT-LLM outperforms all baselines across all datasets and models. On average, its pass@1 exceeds $\mathcal{B}^{4}$ by $13.0\%$, CodeT by $16.6\%$, and Original by $40.7\%$. We assess the statistical significance of these improvements with McNemar’s test (McNemar47), suitable for comparing two algorithms on the same set of tasks with a binary success metric (e.g., pass@1). It confirms that ExPairT-LLM significantly outperforms each baseline on every dataset ($p$-value <0.001<0.001).
@@ Conclusion
Proofs
See 1
Proof.
$C^{*}$ is initially $\mathcal{P}$ (3). We show that, starting from the second iteration, the size of $C^{*}$ decreases by at least one in 12. Thus, in iteration $|\mathcal{P}|$, at the latest, $C^{*}$ contains a single program and ExPairT-LLM terminates. At the end of every iteration, ExPairT-LLM either terminates or finds and validates a differentiating input. In the latter case, the next call to cluster returns at least two clusters which partition $C^{*}$, i.e., their sizes are smaller than the size of $C^{*}$. One of these clusters is set to $C^{*}$ in 12, and thus the claim follows. ∎
See 2
Proof.
Let $p^{*}\in\mathcal{P}$ be a correct program for the task $t$. We prove by induction.
Base: In the first iteration, ${C}^{*}=\mathcal{P}$. Thus, $p^{*}\in{C}^{*}$.
Inductive Hypothesis: Assume that $p^{*}\in{C}^{*}$ at the start of iteration $k$.
Inductive Step: We show that $p^{*}\in{C}^{*}$ at the start of iteration $k+1$. By the inductive hypothesis, $p^{*}\in{C}^{*}$ at the start of iteration $k$. In iteration $k$, the programs in ${C}^{*}$ are clustered by their outputs for the inputs in $\mathcal{I}$. Let $\bar{{C}}$ be the cluster containing $p^{*}$. We show that $\bar{{C}}$ obtains the maximal score in Copeland’s method. Thus, it is set to ${C}^{*}$ (12) and at the start of iteration $k+1$, $p^{*}\in{C}^{*}$. Since the clusters differ by their outputs and since $p^{*}\in\bar{{C}}$, the score of $\bar{{C}}$ is $|C^{*}|-1$. This follows since for every input in $\mathcal{I}$, the output produced by $p^{*}$ (which is identical to the outputs of all programs in $\bar{{C}}$ with respect to $\mathcal{I}$) is more suitable than the outputs of any other cluster. Any other cluster ${C}_{j}\neq\bar{{C}}$ loses a point in the query comparing it to $\bar{{C}}$. Thus, $\bar{{C}}$ has the maximal score. ∎
See 3
Proof.
Assume ${C}^{*}$ contains an incorrect program $p^{\prime}$. By Lemma 2, ${C}^{*}$ contains a correct program $p^{*}$, which is not equivalent to $p^{\prime}$. In 1417, ExPairT-LLM compares the program ${C}^{*}[1]$ with every other program in $C^{*}$. Assume in contradiction that the oracle does not return a differentiating input for any pairwise equivalence query. Namely, it determines that ${C}^{*}[1]$ is equivalent to $p^{*}$ and $p^{\prime}$, implying that $p^{*}$ is equivalent to $p^{\prime}$ – contradiction. ∎
See 4
Proof.
By Lemma 1, ExPairT-LLM terminates. In that iteration, $C^{*}$ contains one program or multiple equivalent programs. This follows by Lemma 3, since in iterations that $C^{*}$ contains nonequivalent programs, the oracle finds a differentiating input, and thus ExPairT-LLM continues to another iteration. When it terminates, ExPairT-LLM returns the first program in $C^{*}$. By Lemma 2, in this iteration, $C^{*}$ contains a correct program. Thus, it returns a correct program. ∎
See 5
Proof.
We define two random variables:
  • $X$: the number of pairwise membership queries in which $C_{n+1}$ wins. Since $C_{n+1}$ is compared once with each cluster, wins with probability $p$, and the queries are independent: $X\sim\mathrm{Binomial}(n,p)$.
  • $Y_{i}$, for $i\in[n]$: the number of pairwise membership queries where $C_{i}$ wins an incorrect cluster. Since $C_{i}$ is compared once with each cluster, wins with probability $0.5$, and the queries are independent: $Y_{i}\sim\mathrm{Binomial}(n-1,0.5)$.
let $B$ be the event that the maximum over all $Y_{i}$ is at most $j$, that is: $\max_{1\leq i\leq n}Y_{i}\,\leq\,j$. By the law of total probability:
$\mathbb{P}(A)=\mathbb{P}(A\;\bigm|B)\cdot\mathbb{P}(B)+\mathbb{P}(A\;\bigm|\overline{B})\cdot\mathbb{P}(\overline{B})$
Since probabilities are nonnegative, we lower bound the second term by 0.
Bounding $\mathbb{P}\bigl(A\mid B\bigr)$: Given $B$, for $C_{n+1}$ to obtain the maximal score, it is sufficient that: $X\;\geq\;j\;+\;2$. Since $X\sim\mathrm{Binomial}(n,p)$, we obtain:
$\mathbb{P}(A\bigm|B)\geq\mathbb{P}\bigl(X\,\geq\,j+2\bigr)$ $\;=\;\sum_{k=\lceil j\rceil+2}^{n}\binom{n}{k}\,p^{k}\,q^{\,n-k}.$
Bounding $\mathbb{P}\bigl(B\bigr)$: By the complement rule, $\mathbb{P}\bigl(B\bigr)=1-\mathbb{P}\bigl(\overline{B}\bigr)$, where $\overline{B}$ is $\max_{1\leq i\leq n}Y_{i}>j$.
By the union bound and since $Y_{i}\sim\mathrm{Binomial}(n-1,0.5)$:
$\mathbb{P}\bigl(\overline{B}\bigr)\;\leq\;\sum_{i=1}^{n}\mathbb{P}\bigl(Y_{i}>j\bigr)\;=\;n\cdot\sum_{k=\lfloor j\rfloor+1}^{\,n-1}\binom{n-1}{k}\Bigl(\tfrac{1}{2}\Bigr)^{n-1}.$
Hence, $\mathbb{P}(B)\;\geq\;1\;-\;n\,\Bigl(\tfrac{1}{2}\Bigr)^{n-1}\sum_{k=\lfloor j\rfloor+1}^{\,n-1}\binom{n-1}{k}$.
Combining all bounds completes the proof. ∎
See 6
Proof.
[…]
$\mathbb{P}(A\,\mid\,B)\;=\;1-\mathbb{P}(\overline{A}\,\mid\,B)=1-(1-p)^{\,d}.$
Similarly,
$\mathbb{P}(A\mid\overline{B})$ $=1-\mathbb{P}(\overline{A}\mid\overline{B})$
$=1-((1-p)^{\lvert{C}^{*}\rvert-d}\cdot(1-p^{\prime})^{d-1})$
Substituting these terms into the law of total probability completes the proof. ∎
See 7
Proof.
We begin with the proof for the pairwise membership queries. We show that every two programs $p_{i},p_{j}\in\mathcal{P}$ can be compared by a pairwise membership query, defined over their containing clusters $C_{i}$ and $C_{j}$, at most once. Thus, the maximal number of pairwise membership queries is $|\mathcal{P}|\choose 2$}. Let $p_{i}$ and $p_{j}$ be programs that are compared by a pairwise membership query in some iteration. Since pairwise membership queries are between different clusters, $p_{i}$ and $p_{j}$ are not in the same cluster. Denote their clusters $C_{i}$ and $C_{j}$, respectively. At the end of this iteration, one cluster is selected as $C^{*}$, and all other clusters (including the programs they contain) are discarded. Thus, at least $p_{i}$ or $p_{j}$ cannot be compared to any other program in future iterations. Next, we show the proof for the pairwise equivalence queries. Consider $C^{*}$ in Line 13. In the first iteration, $|C^{*}|\leq|\mathcal{P}|$ and starting from the second iteration, $|C^{*}|$ decreases by at least one, since the differentiating input leads to splitting the previous $C^{*}$. Since in every iteration ExPairT-LLM poses up to $|C^{*}|-1$ pairwise equivalence queries, it poses at most $\sum_{m=1}^{|\mathcal{P}|-1}m=\frac{|\mathcal{P}|(|\mathcal{P}|-1)}{2}={|\mathcal{P}|\choose 2}$ pairwise equivalence queries.
LLM Prompt Examples for Pairwise Queries
Figure 7 and Figure 8 show prompt examples provided to the LLM oracle for pairwise membership and equivalence queries, respectively. The text in black is the general prompt template, which explains the task and specifies the desired response format. The text in green is the coding task $t$. For the pairwise membership query (Figure 7), the text in red highlights the inputs differentiating the two programs, and the text in blue shows the executed outputs of the two clusters for those inputs. For the pairwise equivalence query (Figure 8), the text in purple is the programs compared for semantic equivalence. If the programs are not equivalent, the LLM is expected to provide a differentiating input.
Example: Pairwise Membership Query You are provided with the partial docstring of a Python function and multiple input-output examples from two different programs (Program 1 and Program 2). Your task is to analyze these examples and determine which program’s outputs better align with the function’s intended behavior as described in the docstring. **Please provide only the classification as a single term: ”Program 1” or ”Program 2”. Do not include any additional text or explanations.** Here is the information provided: - **Docstring:** def max_difference(test_list): """
Write a function to find the maximum difference between available pairs in the given tuple list.
[…]
Output of Program 1:
-1
Output of Program 2:
7
[…]
Output of Program 1:
2
Output of Program 2:
15
Figure 7: An example of a prompt for a pairwise membership query.
Example: Pairwise Equivalence Query You are given the following partial docstring of a Python function: def max_difference(test_list): """
Write a function to find the maximum difference between available pairs in the given tuple list.
and two Python programs implementing this function: **Program 1:** def max_difference(test_list): res = max([(a - b) for a, b in test_list])
return res
**Program 2:** def max_difference(test_list): res = max([abs(b - a) for a, b in test_list])
return res
**Task:** Find an input example such that when this input is provided to both programs, they produce different outputs. **Important:** The input should be **valid according to the function’s description**. **Output Format:** If you find such an input, please return it in the format: parameter_name1=input_value1, parameter_name2=input_value2, … If you cannot find such an input, please return the term ”NO_DIFF”. **Important:** Provide **only** the input in the specified format or the term ”NO_DIFF”, without any additional explanations or output.
Figure 8: An example of a prompt for a pairwise equivalence query.